OpenStudy (anonymous):

PLANS FOR AN ARENA MAP FOR A POPULAR VIDEO GAME ARE DRAWN ON A COORDINATE GRID. A NEW CORRIDOR OF THE ARENA LIES ON LINE Y=5/2x A PERPENDICULAR CORRIDOR PASSES THROUGH THE POINT (10, 0). WRITE THE EQUATION OF THE LINE THAT CONTAINS THE NEW CORRIDOR.

OpenStudy (anonymous):

Y= 5/2x - 25 Y= -2/5x + 4 Y= 2/5x - 4 Y= -5/2x + 25